A non conducting spherical shell of  inner radius "a" and outer radius "b" has p= A/r where a is constant and r is the distance from the center of shell . in addition a positive point charge "q" is locatewd at center . What value should A have if electric field in the shell is to be uniform ?

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

(∫(from r=a to r=s)(A/r)(4pi*r2)dr)/(4πε0s2)=q/(4πε0s2)